I need length and width of triangle.
400=2l+40+2w

Answer by checkley77(12844)

400=2L+40+2W
2L=400-40-2W
L=(400-40-2W)/2
L=(360-2W)/2
L=180-W ANS. FOR THE LENGTH.
2W=400-40-2L
2W=360-2L
W=(360-2L)/2
W=180-L ANS. FOR THE WIDTH.
no REAL SOLUTIONS BECAUSE YOU HAVE 2 UNKNOWNS & ONLY 1 EQUATION.